Topic:gas laws
ques: 500cm^3 of oxygen was collected over water at 30 degree centigrate and 752 mmHg, what is the volume of the dry oxygen at S.T.P(water vapour pressure at 30 degree celcius =3.2mmHg).
1
Expert's answer
2017-05-29T09:57:11-0400
PV=nRT ⇒n=PV/RT
P(O_2 )=P-P(H_2 O)
P(O_2 )=752-3.2=748.8 (mmHg)=998.3 kPa
n=PV/RT= (998.3×0.5)/(8.31×(273+30))=0.2 (mol)
PV=nRT ⇒V=nRT/P
V=(0.2×8.31×273)/101.3=4.44 (〖dm〗^3)

Answer: 4.44 dm3.